OpenStudy (anonymous):

Find the 30th term of the following sequence. 1, 7, 13, 19, ... 174 175 180 181

OpenStudy (anonymous):

try \(1+6\times29\)

OpenStudy (butterflydreamer):

you can use arithmetic progression:  Therefore in your case, if you look at your sequence: 1, 7 , 13, 19 Can you see that that first term is 1? AND the difference between the terms is 6? So basically 1 + 6 = 7 7 + 6 = 13 13 + 6 = 19 , etc. So therefore a = 1 d = 6 and your n (number of terms) = 30 because we're trying to find the 30th term SO,